OpenStudy (anonymous):

Which is the iterative rule for the sequence " 5 , 0 , -5 , -10 , -15" ? an= n+4 an= 5(2-n) an = 10-n an =5 (2-2n)

OpenStudy (anonymous):

PLEEEAS HELP been stuck on this forever

OpenStudy (anonymous):

the answers are above ^^ A B C D

OpenStudy (anonymous):

B

OpenStudy (anonymous):

a1=5(2-1)=5 a2=5(2-2)=0 a3=5(2-3)=-5 a4=5(2-4)=-10 a5=5(2-5)=-15

OpenStudy (anonymous):

so b is the correct answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can you help me answer 1 more please i been stuck on it for hours

OpenStudy (anonymous):

sure